OpenStudy (anonymous):

Simplify cos θ - cos θ sin^2 θ please?

OpenStudy (ybarrap):

use cos^2 + sin^2 = 1

OpenStudy (ybarrap):

distribute out your cos, then you'll have my 1st equation to apply

OpenStudy (anonymous):

\[\cos θ - \cos θ \sin^2 θ = \cos θ(1 - \sin^2 θ)=\cos θ \times \cos^2 θ = \cos^3 \theta\]

OpenStudy (anonymous):

Thank you very much!